OpenStudy (anonymous):

how do i factor this problem... 9-49x^2

OpenStudy (anonymous):

9=3^2 49x^2=(7x)^2

OpenStudy (anonymous):

Now, use a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

Does it help?